OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (anonymous):

What is your x1?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

(2,5) is (x,y) so which is your x1

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

im not the brightest person wen it comes to math. so sorry if i seem a little dumb. im a history person lol

OpenStudy (anonymous):

y-y0=m(x-x0) y-5=3(x-2)

OpenStudy (anonymous):

Yes so the formula for point slope is y-y1=m(x-x1) so plug in the m which is 3

OpenStudy (anonymous):

Then plug in the 5 and the 2 so it would be the last one.